what is the percent discount on a jacket marked down from $120 to $100
Found 3 solutions by rfer, ewatrrr, jgutierrez1975:
Answer by rfer(16322)

20/120=0.17=17%

Answer by ewatrrr(24785)


Hi,
Let x represent the percentage:
Jacket was 20 Dollars off
x/100 *$120 = $20
1.2x = 20
x = 16.67 %
CHECKING our Answer
.1667*$120 = $20

Answer by jgutierrez1975(2)

To find percent decrease:
Subtract the old price (120) by the new price (100) and divide your answer (20) by the old price (120) Always divide by the "old" price or number.
120-100=20
20/120= .17
move decimal two placed to the right and you have 17%